OpenStudy (anonymous):

I need help with exponential functions! I need to figure out when the population will be 1900 using this equation f(t)=650(1.049)^t

jimthompson5910 (jim_thompson5910):

f(t)=650(1.049)^t 1900=650(1.049)^t 1900/650 = (1.049)^t 2.9230769 = (1.049)^t I'll let you finish